OpenStudy (anonymous):

HELP!! PLEASE HELP ME SOLVE THIS! Find the length of the major axis for the ellipse defined by this equation: ((x+3)^2)/(16)+((y-3)^2)/(25)=1

OpenStudy (anonymous):

standard form of this is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] in this case \(a^2=16\implies a=4\) and \(b^2=25\implies b=5\)

OpenStudy (anonymous):

since \(a<b\) you know it looks like this 

OpenStudy (anonymous):

the major axis has length \(2\times 5=10\) and the minor axis has length \(2\times 4=8\)

OpenStudy (anonymous):

Thank you!! This helped a lot! :)